# Model agnostic local variable importance for locally dependent relationships

### v1

@@ Abstract
Global variable importance measures are commonly used to interpret machine learning model results. Local variable importance techniques assess how variables contribute to individual observations rather than the entire dataset. Current methods typically fail to accurately reflect locally dependent relationships between variables and instead focus on marginal importance values. Additionally, they are not natively adapted for multi-class classification problems. We propose a new model-agnostic method for calculating local variable importance, CLIQUE, that captures locally dependent relationships, contains improvements over permutation-based methods, and can be directly applied to multi-class classification problems. Simulated and real-world examples show that CLIQUE emphasizes locally dependent information and properly reduces bias in regions where variables do not affect the response.
@@ 1 Introduction
Variable importance measures are essential for interpreting machine learning models, as they evaluate the relative importance of each feature in the trained model. In particular, global variable importance methods quantify this importance for an entire dataset and have been researched extensively. Breiman [5] introduced a permutation technique within Random Forests, which was later generalized to model-agnostic approaches [9] and extended to better handle collinear features [26, 3]. Variable importance has also been assessed via the leave-one-covariate-out (LOCO) method [16, 13, 4]. The knockoffs technique [2, 6] has also been developed to compute importances while incorporating conditional sampling. Greenwell et al. [11] used partial dependence plots (PDPs) to compute global importance [12].
Local variable importance seeks to measure the relative importance of each feature for individual observations. Existing methods include ICE [10], LIME [24], Counterfactual Explanations [28, 8], Anchors [23], Shapley values [25], and SHAP [18]. SHAP and LIME are particularly popular because of their theoretical foundations, applications to a variety of data types, and their intuitive results [19]. Compared to other local importance techniques, SHAP and LIME stand out due to their yielding an importance value for every value in a dataset. For this reason, we will focus primarily on comparing our results to SHAP and LIME. We demonstrate in our experiments that SHAP and LIME values often yield nonzero importances in regions of the data where a variable should have no impact on the response. This can lead to inaccurate interpretations where locally unimportant variables are falsely characterized as important. Generalizing these methods to multi-class classification tasks is also challenging. This is because they focus on predictions, which requires developing importances for each individual response class.
@@
1.1 Related Work.
Several local variable importance measures currently exist. Shapley values [25] are one of the more popular methods. Shapley values have been extended to SHapley Additive exPlanations or SHAP [18], which ensures that values add up to the prediction of each observation. SHAP values are inspired by game theory in which the goal is to assign payouts to players depending on their contribution to the total payout [27].
SHAP values are calculated by examining coalitions of feature values [19]. This process involves altering the value of a given variable for an observation, while allowing other variables to vary, and observing the effect on the prediction. The average change in prediction across all coalitions yields the SHAP value, capturing not only marginal effects but sometimes interactive ones. However, our experiments indicate that individual SHAP values do not capture local interactions as effectively as CLIQUE.
LIME [24] is another local method which uses interpretable models such as linear regression or classification trees to explain individual predictions of black box models. LIME creates a new dataset of perturbed (noise added) features, where the black box predictions become the response [19]. An interpretable model is built on the new dataset, weighted by proximity to a given observation of interest. The local importances for that observation are then computed from the global variable importances for the interpretable model.
The resulting output requires many simple models: one for each observation. Variables in the models have weights and can be used to create a matrix of weights or importances. We show in our experiments that LIME tends to focus almost entirely on marginal variable importances and fails to detect variable interactions.
In the ICE method, developed by Goldstein et al. [10], a model is built for relating a response to features. Variables are sequentially evaluated by systematically replacing each observation value with values from a uniform grid that ranges from the minimum to maximum value of the variable. While a uniform grid is typical, CLIQUE employs the more robust quantile grid technique. In ICE, predictions are then generated for each grid value. ICE plots are created by plotting the predictions across the grid values in a line graph, where each line represents an observation from the dataset. Molnar [19] points out that ICE plots are inherently local and are often insightful when dealing with variable interactions. However, ICE plots do not natively produce a single measure of variable importance, which can make them difficult to compare directly with CLIQUE.
@@ 2 Local Variable Importances with CLIQUE
Consider a dataset $(\mathbf{x}_{1},y_{1}),\dots,(\mathbf{x}_{n},y_{n})=(X,Y)$, where $X\in\mathbb{R}^{n\times p}$ is the data matrix and $Y\in\mathbb{R}^{n}$ contains the corresponding labels. CLIQUE is model-agnostic and uses cross-validation (CV) predictions to compute local importances. The goal of CLIQUE is to obtain a measure of importance for the $j$th variable at the $i$th data point.
Intuitively, the importance of variable $j$ for the data point $\mathbf{x}_{i}$ in a prediction model can be measured based on how much the model error changes when different values of the variable are used. To achieve this, we first build a full data model, using CV to perform hyper-parameter tuning if necessary. This results in a complete model that could be deployed for prediction on unlabeled data. We then train several CV models with the same hyper-parameters. Each variable is sequentially selected, and its values are altered for each data point. For a given data point, the altered predictions are obtained via the corresponding CV model, which excludes that data point from the training set. The error rate of the CV model with altered values is then compared to the original CV error rate to derive a local importance measure. CV error is used instead of the training error since it more accurately reflects the future performance of the model on test data [19].
Some importance measures can use only a single altered value of a variable [9]. While this approach is generally sufficient for global importance measures that average across data points, it can lead to high variance in local importance measures. To address this, local methods use multiple alterations to reduce variance [18]. Molnar suggests that the most accurate results are achieved by replacing a specific value with all other possible variable values, though this is computationally expensive [19]. We efficiently approximate this by using a quantile grid of values for each variable, spanning the range of variable value.
Mathematically, consider the data point $\mathbf{x}_{i}$ with its corresponding CV model. We aim to compute the local importance for variable $j$, which we denote as $V_{ij}$. Let $\hat{f}(\mathbf{x}_{i})$ represent the original prediction of $y_{i}$ given $\mathbf{x}_{i}$ as input using the CV model where $\mathbf{x}_{i}$ is excluded from training. Let $\tilde{\mathbf{x}}_{i}(j,m)$ be the altered version of $\mathbf{x}_{i}$ in which the $j$th variable is replaced with the $m$th altered value from the quantile grid. The CLIQUE local variable importance is then computed as
(2.1) $V_{ij}=\frac{1}{M}\sum_{m=1}^{M}\mathcal{L}\left(\hat{f}\left(\tilde{\mathbf{x}}_{i}(j,m)\right),y_{i}\right)-\mathcal{L}\left(\hat{f}(\mathbf{x}_{i}),y_{i}\right),$
@@
[…]
3:
4:function CLIQUE($X$, $Y$, model, $M$)
5:     $mod\leftarrow$ model($X$, $Y$)
6:     $cvmod\leftarrow$ individual CV models
7:     $\underset{n\times 1}{Err}$leftarrow $\mathcal{L}$(predict($cvmod$, $X$), $Y$)
8:     $\underset{n\times p}{\mathrm{V}}\leftarrow\underset{n\times p}{0}$
9:     for $j$ in $\underset{p\times 1}{vars}$ do
10:         $\underset{M\times 1}{grid}\leftarrow$ $M$ quantile values for variable $j$
11:         for $m$ in $\underset{M\times 1}{grid}$ do
12:              $\underset{n\times p}{W}\leftarrow\underset{n\times p}{X}$
13:              $\underset{n\times 1}{W_{:j}}\leftarrow m$
[…]
15:              $\underset{n\times 1}{\mathrm{V}_{:j}}\leftarrow\mathrm{V}_{:j}+\frac{\widetilde{Err}_{j}-Err}{M}$
16:         end for
17:     end for
18:     return $\underset{n\times p}{\mathrm{V}}$
19:end function
We note that CLIQUE captures conditional information from the perspective that a conditional effect is present if it depends or changes upon another effect. Thus, we are not referring to conditional information in a formal probabilistic sense. We also note that for our work, “local” suggests at least some conditionality or location-dependence on the importance of a variable. If there is no conditionality, then local importance should be mostly redundant with global importance, which can be assessed using existing global methods.
Some global importance approaches permute variable values instead of using a grid [5, 9]. To demonstrate that using the quantile grid in CLIQUE outperforms repeated permutations, we include in our experimental results a method we refer to as Conditional Local Importance by Permutations (CLIP). CLIP and CLIQUE are nearly identical, except that CLIP performs $M$ variable permutations to select the alternative values, while CLIQUE replaces a variable with $M$ quantile values.
@@ 3 Simulated Experiments
To demonstrate how CLIQUE outperforms the popular methods SHAP and LIME in capturing locally dependent relationships, we compare the methods on simulated data with known structures. For all analyses, we use $M=25$ in both CLIP and CLIQUE.
For each of the following three data simulations, we let all features be independent and identically distributed (iid) from a uniform distribution $\mathcal{U}(-1,1)$. Here, iid implies that all of the features are orthogonal to each other. We generate 400 training observations for each dataset and then produce the response variable values using a simulation specific equation.
3.1 AND Gate Data.
@@
We begin with a simple simulation that uses a classic concept from digital and mathematical logic known as an AND gate (Figure 1). To create this data, we generate three features and produce the response using:
(3.2) $y=\begin{cases}1,&\text{if }v_{1}>-\frac{1}{3}\text{ \& }v_{2}>-\frac{1}{3}\\ 0,&\text{otherwise.}\end{cases}$
Note that the third feature, $v_{3}$ is simply a noise variable that has no influence on the output.
Figure 1: Scatterplot of the AND gate data colored by the label $y$ as computed in Equation 3.2. When $v_{2}<-1/3$, the value of $v_{1}$ should have no impact on the output of the trained model and should thus have no importance.
@@
We then build a Random Forest model using the randomForest package [17] in the R programming language [22]. From this model, we extract SHAP values using the treeshap package [15] and LIME values using the lime package [14]. We compare their values with CLIP and CLIQUE values in Figure 2.
@@
Figure 2: Scatterplots of local variable importances vs. $v_{1}$ values from the AND gate data experiment. Each plot is colored by whether $v_{2}>-1/3$. CLIQUE and CLIP values match our expectations by outputting an importance of zero when $v_{2}<-1/3$ and a nonzero importance otherwise. LIME values fail to distinguish between these two regions and SHAP values are nonzero when $v_{2}<-1/3$.
@@
Based on the geometry of the data in Figure 1, we can establish several expectations for the local variable importances. When $v_{2}<-1/3$, the values of $v_{1}$ should have no impact on the model output. Therefore, we expect the importance for $v_{1}$ to be close to zero when $v_{2}<-1/3$. However, the values of $v_{1}$ should have large positive importance measures when $v_{2}>-1/3$ due to their significant impact on the model output.
@@
From Figure 2, we see this in the CLIQUE values. When $v_{2}<-1/3$, the importance of $v_{1}$ is practically zero, while for $v_{2}>-1/3$, $v_{1}$ has a positive importance. CLIP shows a similar trend, though with higher variance in the nonzero importances, highlighting the advantage of using the quantile grid in CLIQUE instead of permutations. In contrast, LIME shows no discernible difference in the importance of $v_{1}$ relative to $v_{2}$, indicating that it fails to capture the conditional information in the data. SHAP falls between LIME and CLIQUE. It displays two distinct trends for $v_{1}$, but still assigns nonzero importance when $v_{2}<-1/3$. This could mislead users to believe that $v_{1}$ is locally important when it should not have any impact on the prediction.
3.2 Corners Data.
@@
Figure 3: Scatterplot of the Corners data colored by the label $y$ (see Equation 3.3). $v_{2}$ is not important when $v_{1}<0$. $v_{1}$ is not important when $|v_{2}|<1/4$.
@@
The AND Gate is a symmetric problem, where the results for $v_{1}$ mirror the results for $v_{2}$. We now simulate values for a non-symmetric dataset that we call the Corners data (Figure 3). We generate three features and produce the response using:
(3.3) $y=\begin{cases}1,&\text{if }v_{1}>0\text{ \& }|v_{2}|>\frac{1}{4}\\ 0,&\text{otherwise}.\end{cases}$
@@
In this data, we expect that $v_{2}$ is not important when $v_{1}<0$ and $v_{1}$ is not important when $|v_{2}|<1/4$. We use Generalized Boosted Modeling (GBM) on the Corners data and extract LIME, SHAP, CLIP, and CLIQUE values, which are shown in Figures 4 and 5. Again the CLIQUE and CLIP values reflect the expected behavior, with CLIQUE showing a smaller variance than CLIP. In contrast, the LIME values again show only the marginal information. SHAP values primarily highlight the marginal effects with only a small amount of conditional information.
3.3 Regression Interaction Data.
@@
(3.4) $y=\begin{cases}v_{1},&\text{if }v_{3}>0\\ v_{2},&\text{if }v_{3}<0.\end{cases}$
In this data, $v_{1}$ should not be important when $v_{3}<0$ and $v_{2}$ should not be important when $v_{3}>0$. We build a Random Forest model and again extract LIME, SHAP, CLIP and CLIQUE values, with the results shown in Figure 7. The trends are similar to previous findings: CLIP and CLIQUE align with the expected relationships, while LIME fails to do so. As in Section 3.1, SHAP values show differences between the two regions, but most values remain nonzero for the region where no $v_{1}$ importance is expected. A unique observation for this regression problem is that CLIP and CLIQUE show a quadratic relationship with data values when they are locally important.
Figure 6: Scatterplot of the Regression Interaction data colored by whether $v_{3}>0$ (see Equation 3.4). $v_{1}$ is not important when $v_{3}<0$. $v_{2}$ is not important when $v_{3}>0$.

### v2

@@ Abstract
Global variable importance measures are commonly used to interpret the results of machine learning models. Local variable importance techniques assess how variables contribute to individual observations. Current methods typically fail to accurately reflect locally dependent relationships between variables and instead focus on marginal importance values. Additionally, they are not natively adapted for multi-class classification problems. We propose a new model-agnostic method for calculating local variable importance, CLIQUE, that captures locally dependent relationships, improves over permutation-based methods, and can be directly applied to multi-category classification problems. Simulated and real-world examples show that CLIQUE emphasizes locally dependent information and properly reduces bias in regions where variables do not affect the response.
@@ 1 Introduction
Variable importance measures are essential for interpreting machine learning models, as they evaluate the relative importance of each feature in the trained model. In particular, global variable importance methods quantify this importance for an entire dataset and have been researched extensively. Breiman [4] introduced a global permutation technique for an internal validation set within Random Forests. This concept was later generalized to model-agnostic approaches [8]. Other global variable importance methods include leave-one-covariate-out (LOCO) [15, 12, 3], knockoffs [2, 5], and using the spread of predictions in partial dependence plots (PDPs) [10, 11]
Local variable importance seeks to measure the relative importance of each feature for individual observations. Existing methods include ICE [9], LIME [24], Counterfactual Explanations [27, 7], Anchors [25], Shapley values [26], and SHAP [17]. SHAP and LIME are particularly popular because of their theoretical foundations, applications to a variety of data types, and their intuitive results [19]. SHAP values are calculated by considering coalitions of feature values. This involves changing the value of a given variable for a given observation, while allowing other variables to change, and observing the impact on prediction. The average change in prediction across all coalitions yields the SHAP value. Because this allows other variables to change, SHAP values can capture effects that are not only marginal, but sometimes even interactive. However, our experiments show that individual SHAP values do not always capture local interactions effectively.
LIME [24] is another local method which uses interpretable models such as linear regression or classification trees to explain individual predictions of black-box models. LIME creates a new dataset of perturbed features (noise added), where the black-box predictions become the response [19]. An interpretable model is built on the new dataset, weighted by proximity to a given observation of interest. The local importances for that observation are then computed from the global variable importances for the interpretable model. The resulting output requires many simple models: one for each observation. The variables in the models have weights and can be used to create a matrix of weights or importances. We show in our experiments that LIME tends to focus almost entirely on marginal variable importances and fails to detect variable interactions.
In the ICE method, developed by Goldstein et al. [9], a model is built to relate a response to features. Variables are sequentially evaluated by systematically replacing each observation value with values from a uniform grid that ranges from the minimum to the maximum value of the variable. Although a uniform grid is typical, CLIQUE employs the more robust quantile grid technique. In ICE, predictions are then generated for each grid value. ICE plots are created by plotting the predictions across the grid values in a line graph, where each line represents an observation from the dataset. Molnar [19] points out that ICE plots are inherently local and are often insightful when dealing with variable interactions. However, ICE plots do not natively produce a single measure of variable importance.
Compared to other local importance techniques, SHAP and LIME stand out because they produce an importance value for every value in a data set. For this reason, we will focus primarily on comparing our results with those of SHAP and LIME. We demonstrate in our experiments that SHAP and LIME values often yield nonzero importances in regions of the data where a variable should have no impact on the response. This can lead to inaccurate interpretations where locally unimportant variables are falsely characterized as important. Generalizing these methods to multi-category classification tasks is also challenging. This is because they focus on predictions, which requires developing importances for each individual response class.
@@ 2 Local Variable Importances with CLIQUE
Consider a dataset $(\mathbf{x}_{1},y_{1}),\dots,(\mathbf{x}_{n},y_{n})=(X,Y)$, where $X\in\mathbb{R}^{n\times p}$ is the data matrix and $Y\in\mathbb{R}^{n}$ contains the labels. CLIQUE is model-agnostic and uses cross-validation (CV) predictions to compute local importances. The goal of CLIQUE is to obtain a measure of importance for the $j$th variable at the $i$th data point. Intuitively, the importance of variable $j$ for a given data point $\mathbf{x}_{i}$ in a prediction model can be measured by assessing how much the model error changes when the variable values change.
CLIQUE aims to achieve several desirable properties when assessing local importance:
  1. P1
    If altering a variable value does not affect the model output for an observation, that variable should have an importance near 0 for that observation.
  2. P2
    An immediate application to multi-class problems, where if altering a variable yields new predicted classes or probabilities, that variable is important regardless of the class label it has.
  3. P3
    Importance directly depends on model errors rather than predictions.
To achieve a method that offers these properties, we first build a full model using CV for hyperparameter tuning. Although training errors can be used to assess importance, test errors provide a better estimate of model performance on unseen data [19]. To capitalize on these benefits, CLIQUE computes local importance using CV errors by training multiple CV models with the same hyperparameters as the final model. Each variable is systematically modified for every data point and predictions are generated from a CV model that excluded that point from training. Comparing the CV error rate of the modified data to the original CV error rate yields a local importance measure.
Some importance measures can use only a single altered value of a variable [8]. While this approach is generally sufficient for global importance measures that average across data points, it can lead to high variance in local importance measures. To address this, local methods use multiple alterations to reduce variance [17]. Molnar suggests that the most accurate results are achieved by replacing a specific value with all other possible variable values, although this is computationally expensive [19]. We efficiently approximate this by using a quantile grid of values for each variable, spanning the variable range.
Mathematically, consider the data point $\mathbf{x}_{i}$ with its corresponding CV model. We aim to compute the local importance for variable $j$, which we denote as $V_{ij}$. Let $\hat{f}(\mathbf{x}_{i})$ represent the original prediction of $y_{i}$ given $\mathbf{x}_{i}$ as input using the CV model where $\mathbf{x}_{i}$ is excluded from training. Let $\tilde{\mathbf{x}}_{i}(j,m)$ be the altered version of $\mathbf{x}_{i}$ in which the $j$th variable is replaced with the $m$th altered value from the quantile grid. The CLIQUE value is
(1) $V_{ij}=\frac{1}{M}\sum_{m=1}^{M}\mathcal{L}\left(\hat{f}\left(\tilde{\mathbf{x}}_{i}(j,m)\right),y_{i}\right)-\mathcal{L}\left(\hat{f}(\mathbf{x}_{i}),y_{i}\right),$
@@
1:  Input: data $\underset{n\times p}{X}$, labels $\underset{n\times 1}{Y}$, model, quantiles $M$
2:  $mod\leftarrow$ model($X$, $Y$)
3:  $cvmod\leftarrow$ individual CV models
4:  $\underset{n\times 1}{Err}$leftarrow $\mathcal{L}$(predict($cvmod$, $X$), $Y$)
5:  $\underset{n\times p}{\mathrm{V}}\leftarrow\underset{n\times p}{0}$
6:  for $j=1$ to $p$ do
7:     $\underset{M\times 1}{grid}\leftarrow$ $M$ quantile values for variable $j$
8:     for $m=1$ to $M$ do
9:        $\underset{n\times p}{W}\leftarrow\underset{n\times p}{X}$
10:        $\underset{n\times 1}{W_{:j}}\leftarrow m$
@@
12:        $\underset{n\times 1}{\mathrm{V}_{:j}}\leftarrow\mathrm{V}_{:j}+\frac{\widetilde{Err}_{j}-Err}{M}$
13:     end for
14:  end for
We note that CLIQUE captures conditional information from the perspective that a conditional effect is present if it depends or changes upon another effect. Thus, we are not referring to conditional information in a formal probabilistic sense. We also note that for our work, “local” suggests at least some location-dependence on the importance of a variable. If there is no location-dependence, then local importance should be mostly redundant with global importance, which can be assessed using existing global methods.
Many global importance approaches permute variable values instead of using a grid [4, 8]. To show that using the quantile grid in CLIQUE outperforms repeated permutations, we include in our experimental results a method we refer to as Conditional Local Importance by Permutations (CLIP). CLIP and CLIQUE are nearly identical, except that CLIP performs $M$ variable permutations to select alternative values, while CLIQUE replaces a variable with $M$ quantile values.
@@ 3 Simulated Experiments
To demonstrate how CLIQUE outperforms the popular methods SHAP and LIME in capturing locally dependent relationships, we compare the methods on simulated data with known structures. For all analyses, we use $M=25$ in both CLIP and CLIQUE. $M=25$ was initially chosen based on desired qualities of the Central Limit Theorem for stabilizing estimates and reducing variance. Further exploratory analysis of the impact of $M$ on CLIQUE values and their variance is available in Appendix A.3. For each of the following three data simulations, all features are independent and identically distributed (iid) from a uniform distribution $\mathcal{U}(-1,1)$, and we generate 400 training observations.
3.1 AND Gate Data
@@
We begin with a simple simulation based on an AND gate (Figure 1). To create this data, we generate three features and produce the response using:
(2) $y=\begin{cases}1,&\text{if }v_{1}>-\frac{1}{3}\text{ \& }v_{2}>-\frac{1}{3}\\ 0,&\text{otherwise.}\end{cases}$
Note that $v_{3}$ has no influence on the output.
@@
We build a Random Forest model using the randomForest package [16] in the R programming language [22]. From this model, we extract SHAP values using the treeshap package [14] and also LIME values [13].
Figure 1: Scatterplot of the AND gate data colored by the label $y$ as computed in Equation 2. When $v_{2}<-1/3$, the value of $v_{1}$ should have no impact on the output of the trained model and should thus have no importance.
@@
Figure 2: Scatterplots of local variable importances vs. $v_{1}$ values from the AND gate data experiment. Each plot is colored by whether $v_{2}>-1/3$. CLIQUE and CLIP values output an importance of zero when $v_{2}<-1/3$ and a nonzero importance otherwise. LIME values fail to distinguish between these two regions and SHAP values are nonzero when $v_{2}<-1/3$.
@@
Based on the geometry of the data in Figure 1, we expect several properties for the local variable importances. When $v_{2}<-1/3$, the values of $v_{1}$ should have no impact on the model output. Therefore, the $v_{1}$ importance should be close to zero when $v_{2}<-1/3$. However, the values of $v_{1}$ should have large positive importance measures when $v_{2}>-1/3$ due to their significant impact on the model output.
@@
From Figure 2, we see this in the CLIQUE values. When $v_{2}<-1/3$, the importance of $v_{1}$ is practically zero, while for $v_{2}>-1/3$, $v_{1}$ has a positive importance. CLIP shows a similar trend, though with higher variance in the nonzero importances, highlighting the advantage of using the quantile grid in CLIQUE instead of permutations. In contrast, LIME shows no discernible difference in the importance of $v_{1}$ relative to $v_{2}$, indicating that it fails to capture the conditional information in the data. SHAP falls between LIME and CLIQUE. It displays two distinct trends for $v_{1}$, but still assigns nonzero importance when $v_{2}<-1/3$. In other words, CLIQUE satisfies property P1 while LIME and SHAP do not.
3.2 Corners Data
@@
Figure 3: Scatterplot of the Corners data colored by the label $y$ (see Equation 3). $v_{2}$ is not important when $v_{1}<0$. $v_{1}$ is not important when $|v_{2}|<1/4$.
@@
The AND Gate is a symmetric problem, where the results for $v_{1}$ mirror the results for $v_{2}$. We now simulate values for a non-symmetric dataset that we call the Corners data (Figure 3). We generate three features with response:
(3) $y=\begin{cases}1,&\text{if }v_{1}>0\text{ \& }|v_{2}|>\frac{1}{4}\\ 0,&\text{otherwise}.\end{cases}$
In this data, we expect that $v_{2}$ is not important when $v_{1}<0$ and $v_{1}$ is not important when $|v_{2}|<1/4$. To illustrate the model-agnostic capability of CLIQUE, we use Generalized Boosted Modeling (GBM) on the Corners data and extract LIME, SHAP, CLIP, and CLIQUE values, which are shown in Figures 4 and 5. Again, the CLIQUE and CLIP values reflect the expected behavior, with CLIQUE showing a smaller variance than CLIP. In contrast, the LIME values again show only marginal information. SHAP values primarily highlight the marginal effects with only a small amount of conditional information. That is, CLIQUE satisfies property P1 while LIME and SHAP do not.
@@
3.3 Regression Interaction Data
@@
(4) $y=\begin{cases}v_{1},&\text{if }v_{3}>0\\ v_{2},&\text{if }v_{3}<0.\end{cases}$
Figure 6: Scatterplot of the Regression Interaction data colored by whether $v_{3}>0$ (see Equation 4). $v_{1}$ is not important when $v_{3}<0$. $v_{2}$ is not important when $v_{3}>0$.
@@
In this data, $v_{1}$ should not be important when $v_{3}<0$ and $v_{2}$ should not be important when $v_{3}>0$. We build a Random Forest model and again extract LIME, SHAP, CLIP and CLIQUE values, with the results shown in Figure 7. The trends are similar to previous findings: CLIQUE satisfies P1 while LIME and SHAP fail to do so. As in Section 3.1, SHAP values show differences between the two regions, but most values remain nonzero for the region where no $v_{1}$ importance is expected. A unique observation for this regression problem is that CLIP and CLIQUE show a quadratic relationship with data values when they are locally important.